Determine the mass in grams of Li+ (6.941 g/mol) in 25.0 g of Li2SO4 (109.94 g/mol)


1
Expert's answer
2021-03-02T01:59:22-0500

The compound dissociates as the following:

Li2SO4 = 2Li+ + SO42-

From here, the mass of lithium ions equals:

m(Li+) = m(Li2SO4) × 2 × Mr(Li+) / Mr(Li2SO4) = 25.0 g × 2 × 6.941 g/mol / 109.94 g/mol = 3.16 g


Answer: 3.16 g